OpenStudy (anonymous):

Integrate 2/(root 4 - e^(2t) )

OpenStudy (dumbcow):

i think you use the sub: e^t = 2sin(theta) but im verifying that it works

OpenStudy (anonymous):

or e^t=u ?

OpenStudy (dumbcow):

ok it will work yes that sub will work too but eventually you have to use a trig substitution

OpenStudy (anonymous):



OpenStudy (anonymous):

should we take s=4-u^2 ?

OpenStudy (anonymous):

no...

OpenStudy (dumbcow):

wait u = e^t du = e^t dt \[\rightarrow \int\limits_{?}^{?} \frac{2}{u \sqrt{4-u^{2}}} du\]

OpenStudy (anonymous):

ok..makes sense

OpenStudy (dumbcow):

then \[u = 2\sin \theta\] \[du = 2 \cos \theta d \theta \] \[\rightarrow \int\limits_{}^{}\frac{2}{(2\sin \theta)\sqrt{4-4\sin^{2} \theta}} (2\cos \theta) d \theta\] simplify

OpenStudy (anonymous):

1/sin theta !

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

which means, we integrate sin^-1 theta....damn

OpenStudy (dumbcow):

this will help http://math2.org/math/integrals/more/csc.htm